A stone is thrown straight downward with initial speed 9.4 m/s from a height of 29m. Find the time it takes to reach the ground.
1
Expert's answer
2021-10-26T17:18:13-0400

"v_0 = 9.4 \;m/s \\

y_0 = 29 \;m \\

y=y_0+v_0t+ \frac{1}{2}gt^2 \\

0 = 29 +9.4t+\frac{1}{2}(-9.8)t^2 \\

4.9t^2 -9.4t- 29 =0 \\

t = \frac{9.4 + \sqrt{9.4^2 + 4(4.9)(29)}}{2(4.9)} \\

t = 3.57 \;s"

Take note that we only consider the positive value of time.